OpenStudy (anonymous):

find the derivative of y with respect to x in y=sin^(-2)(x^3)

OpenStudy (anonymous):

-2Sin(x^2) * 2x

OpenStudy (anonymous):

* Cos (x^2) <= forgot that part

OpenStudy (anonymous):

Answer is -6x^2cosx^3/ sin^3x^3

OpenStudy (anonymous):

is the (-2)(x^3) all in the exponent?

OpenStudy (anonymous):

\[Sin ^{-2}x ^{2}\]??

OpenStudy (anonymous):

\[\sin^{-2} x^3\]

OpenStudy (anonymous):

Answer is \[-6x^2cosx^3/\sin^3x^3\]

OpenStudy (anonymous):

i think that there should be an x after the sin?!?!?!?

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

wierd! hold on!

OpenStudy (anonymous):

derivative of sin^(-2)(x^3) = -2 * 3x² sin^(-3) x³ * cos x³ = - 6x² cos x³ / sin³ x³